OpenStudy (firejay5):

Show work/explain work to get the Medal for patronage! :D A painter leans a 15-ft ladder against a house. The base of the ladder is 5 ft from the house. a. To the nearest tenth of a foot, how high on the house does the ladder reach?

OpenStudy (firejay5):

@jim_thompson5910

OpenStudy (anonymous):

c^2=A^2+b^2 15^2-5^2=A^2 A^2=sqrt(200) A=14.1 It's just Pythagorean theorem.

OpenStudy (whpalmer4):

 by Pythagorean theorem, \[5^2+x^2=15^2\]Solving for \(x\) \[x^2=15^2-5^2 = 225-25 = 200\]Take square root of both sides\[x = \sqrt{200} = \sqrt{100*2} = 10\sqrt{2}\approx 14.14213562\]so to the nearest 10th of a foot, the answer is 14.1 feet.

OpenStudy (whpalmer4):

drawing a picture and labeling it will often get you quite a ways down the road to the solution.

OpenStudy (firejay5):

I know that, I would've done that if I know which was the legs and hypotenuse and now I know